40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 5 20:37:34 2011, jamie, Summary, General, Summary of today's in-vacuum work 
    Reply  Sat Aug 6 03:10:05 2011, Suresh, Summary, General, Summary of today's in-vacuum work ETMX_OSEMs.pngAfter_osem_adjustment.JPGBefore_osem_adj.JPG
       Reply  Sat Aug 6 13:38:02 2011, rana, Summary, General, Summary of today's in-vacuum work 
          Reply  Sun Aug 7 11:10:54 2011, Suresh, Summary, General, OSEM positioning 
    Reply  Sun Aug 7 14:01:58 2011, kiwamu, Update, Green Locking, X green beam re-aligned 
       Reply  Sun Aug 7 18:41:12 2011, Jenne, Update, Green Locking, X green beam re-aligned 
Message ID: 5127     Entry time: Fri Aug 5 20:37:34 2011     Reply to this: 5130   5133
Author: jamie 
Type: Summary 
Category: General 
Subject: Summary of today's in-vacuum work 

[Jamie, Suresh, Jenne, Koji, Kiwamu]

After this morning's hiccup with the east end crane, we decided to go ahead with work on ETMX.

Took pictures of the OSEM assemblies, we laid down rails to mark expected new position of the suspension base.

Removed two steering mirrors and a windmill that were on the table but where not being used at all.

Clamped the test mass and moved the suspension to the edge of the table so that we could more easily work on repositioning the OSEMs.  Then leveled the table and released the TM.

Rotated each OSEM so that the parallel LED/PD holder plates were oriented in the vertical direction.  We did this in the hopes that this orientation would minimize SD -> POS coupling.

For each OSEM, we moved it through it's full range, as read out by the C1:SUS-ETMX_{UL,UR,LL,LR,SD}PDMon channels, and attempted to adjust the positions so that the read out was in the center of the range (the measured ranges, mid values, and ultimate positions will be noted in a follow-up post).  Once we were satisfied that all the OSEMs were in good positions, we photographed them all (pictures also to follow).

Re-clamped the TM and moved it into it's final position, using the rails as reference and a ruler to measure as precisely as possible :

ETMX position change: -0.2056 m = -20.56 cm = -8.09 in (away from vertex)

Rebalanced the table.

Repositioned the mirror for the ETMX face camera.

Released TM clamps.

Rechecked OSEM centering.

Unblocked the green beam, only to find that it was displaced horizontally on the test mass about half an inch to the west (-y).  Koji determined that this was because the green beam is incident on the TM at an angle due to the TM wedge.  This presents a problem, since the green beam can no longer be used as a reference for the arm cavity.  After some discussion we decided to go with the TM position as is, and to realign the green beam to the new position and relock the green beam to the new cavity.  We should be able to use the spot position of the green beam exiting the vacuum at the PSL table as the new reference.  If the green X beam exiting at the PSL table is severely displaced, we may decide to go back in and move ETMX to tweak the cavity alignment.

At this point we decided that we were done for the day.  Before closing up, we put a piece of foil with a hole in it in front of the the TM face, to use as an alignment aperture when Kiwamu does the green alignment.

Kiwamu will work on the green alignment over the weekend.  Assuming everything works out, we'll try the same procedure on ETMY on Monday.

ELOG V3.1.3-